Given two complex numbers "z" and "w" such that "z=(1+i)w+(3-i)\overline{w}" , where "\bar{w}" is the conjugate of "w" . Deduce what is "\overline{w}" in terms of "z" and "\overline{z}"


1
Expert's answer
2021-10-21T14:31:07-0400

z=(1+I)w+(3-i) conjugate (w)

"z=(1+i)(3-i)\bar{w}...................(1)"


"\\\bar{z}=\overline{(1+i)w+(3-i)\bar{w}}\\"


"=\overline{(1+i)}.\overline{w}+\overline{(3-i)}.\overline{\bar{w}}\\"


"=(1-i).\overline{w}+(3+i)w........................(2)"


Multiplying (3+i) and (1+i) in equation (1) and (2), we get;

"z=(1+i)w+(3-i)\bar{w}\times(3+i)\\\bar{z}=(1-i)\bar{w}+(3+i)w\times(1+i)"


"(3+i)z=(1+i)(3+i)w+(3-i)(3+i)\bar{w}............(iii)\\(1+i)\bar{z}=(1-i)(1+i)\bar{w}+(3+i)(1+i)w............(iv)"


subtracting equation (4) from equation (3) we get


"\bar{w}[-(1+1)+(9+1)]=z(3+i)-(1+i)\bar{z}"


"\bar{w}=\frac{1}{8}[(3+i)z-(1+i)\bar{z}"